OpenStudy (anonymous):

MEDAL AND FANS!!! :) What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (anonymous):

evidently it is \(x-y=\text{something}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

check an see hwat is \(2-7\) ?

OpenStudy (anonymous):

-5 right

OpenStudy (haseeb96):

m=y2-y1 \x2-x1 here (x1,y1) = (-2,3) (x2,y2)=(2,7) so slope is m = 3-7 / 2- (-2) = -4 / 2 +2 = -4 /4 m = -1 is the slope of the line and now take any point and this m=-1 here the point is (-2,3) and put it in this formula y-y1 = m(x-x1) y-3 = -1 (x-3) y-3 =-x +3 y+x -3 -3 =0 x+y -6 =0 is the equation of the line passing from the points .

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

\(-2-3=-5\) pick \(x-y=-5\)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

yw